OpenStudy (anonymous):

What is the output of f(x)=-2-3x when x=-4/3 I think there is something missing i couldn't figure it out.

OpenStudy (anonymous):

Nothing is missing. Just plug in x for -4/3. -2-3(-4/3) = 4-2= 2 You are just solving for f(x). That is really what the question wants.

OpenStudy (anonymous):

Remember: f(x) means function of x, which is just a fancy way of saying y.

OpenStudy (anonymous):

ok i just never realized that. Thanks for the help

OpenStudy (anonymous):

You're welcome!